# Counterfactual Token Generation in Large Language Models

### v1

@@ 4.2 How similar is counterfactually generated text to the factual one?
Experimental setup. We first use the implementation of our model on Llama 3 8B-instruct to generate (factual) outputs to $2{,}000$ question prompts sourced from the LMSYS Chat 1M dataset [51]. As a system prompt we use “Keep your replies short and to the point.”. Further, for each factual output, we perform two interventions where we replace a randomly selected token $t_{i}$ with a token $t^{\prime}\neq t_{i}$. [^7] To select $t^{\prime}$, we set the probability of $t_{i}$ in $d_{i}$ to 00, re-scale the values of $d_{i}$ and use top-$p$ sampling with $p=0.9$. One of the two interventions restricts the choice of $t_{i}$ to the first half of the output sequence and the other restricts it to the second half. Then, for each intervened factual output, we feed the concatenation of the question prompt and the first part of the intervened factual output up to and including token $t^{\prime}$ as input to our model. We regenerate the second part of the output after token $t^{\prime}$ using (i) interventional token generation and (ii) counterfactual token generation, as described in Section 4.1. Finally, we measure the lexicographic similarity between the regenerated second part of the output and its factual counterpart using their (normalized) Levenshtein edit distance [52]. In our experiments, we implement our model using the Gumbel-Max SCM defined in Equation 3 as well as the top-$p$ Gumbel-Max SCM and top-$k$ Gumbel-Max SCM discussed at the end of Section 3.
Results. Figure 3 summarizes the results, which show that the output sequences generated using counterfactual token generation are more similar to the factual sequences (i.e., the edit distance is lower) than the output sequences generated using interventional token generation. This suggests that, even though the top-$p$ and top-$k$ Gumbel-Max SCMs are not guaranteed to satisfy counterfactual stability, in practice, counterfactual token generation under both models does “prioritize” selecting the same tokens $T_{i}$ that were selected during the factual generation.
@@ 4.3 Does counterfactual token generation reveal model biases?
Common approaches to addressing questions of bias and fairness rely on making counterfactual comparisons based on sensitive attributes [53]. For example, would a person’s income have been the same if their race or sex were different? In this section, we focus on a census data generation task, and demonstrate the use of counterfactual token generation to investigate potential biases of the LLM towards demographic groups.
Experimental setup. We first use the implementation of our model on Llama 3 8B-instruct to generate (factual) census data. To this end, we use the same input prompt three times with different seeds (see Appendix B for details), requesting $50$ individuals each time. The factual data generated by the model consist of $114$ fictional individuals including their name, age, sex, citizenship, race, ethnicity, marital status, number of children, occupation, income and education, in this given order. For each fictional person, we consider all possible interventions on each of the sensitive attributes of sex and race. Then, for each intervention, we concatenate the input prompt with the initial part of the output that includes the fictional person’s description (up to and including the intervened sensitive attribute). This concatenated input is then used by our model to regenerate the latter part of the output, following the intervention, using counterfactual token generation (i.e., Algorithm 1). Finally, we compare the factual and counterfactual values of attributes such as income, education and occupation.
(a) Change in income
upon intervention on sex
(b) Change in education level
@@
Figure 4: Comparison between factual and counterfactual income, education, and occupation. Panel (a) shows the income of male (female) individuals had they been female (male). Enlarged points correspond to the median income. Panel (b) shows the average difference in the education level of individuals of each race had their race been different. Here, positive values indicate an improvement in education, negative values indicate a decline, and each race is represented with a short description for visibility; refer to Appendix B for further details. Panel (c) shows the distribution shift of occupations among Asian American individuals had they been Black or African American. Green (red) sections indicate the increase (decrease) in the number of Asian American individuals that practice each occupation had they been Black or African American. In all experiments, the temperature parameter is set to $\tau=0.8$.
@@
Results. Figure 4 summarizes the results, which reveal several interesting insights. Figure 4(a) shows that, for most male individuals, their generated income would have decreased had they been female, whereas, for female individuals, it would have sometimes increased and sometimes decreased had they been male. This suggests that the model of the world constructed by the LLM does not only present bias but also exhibits inconsistencies in its perceived relationship between a person’s sex and income. Figure 4(b) shows that, for individuals of all (generated) races, there exists at least one other race that, had they belonged to it, they would have experienced a significant increase or decrease in their education level (refer to Appendix B for the assignment of each education level to a numerical value). Finally, Figure 4(c) shows that, for Asian American individuals, their occupation would have shifted from STEM to humanities-related occupations had they been Black or African American.
@@ 5 Conclusions
Our work opens many interesting avenues for future work. Our causal model of the autoregressive process underpinning large language models crucially relies on the Gumbel-Max SCM. However, it would be interesting to understand the sensitivity of counterfactual token generation to the specific choice of SCM, and to consider alternative SCMs, especially those that do not satisfy the property of counterfactual stability. Furthermore, we have showcased our model on a single LLM, namely Llama 3 8B-instruct. It would be useful to implement our model on other LLMs and use counterfactual token generation to reveal similarities and differences between the underlying models of the world constructed by different LLMs. Specifically, it would be insightful to see whether the sensitivity of an LLM’s counterfactual output changes as its number of parameters increases. Lastly, an interesting future direction would be to explore the use of our methodology in conjunction with human feedback to train (or fine-tune) LLMs that better understand causal relationships.
@@ Appendix A Additional counterfactual stories
In this section, we provide additional details about the census generation experiment discussed in Section 4.3. Figure 6 shows the complete system and user prompts used to generate the census data. For race and ethnicity, we instructed our model, through the system prompt, to select values among those reported in the latest (2020) US Census. We used this prompt three times with different seeds. Despite our request for $50$ individuals per generation, the LLM only generated $34$, $39$ and $41$ individuals each time, resulting in a total of $114$ individuals. Table 1 contains the full descriptions of the race attribute values, of which shortened versions were used in Figure 4(b). Finally, Table 2 lists the numerical values assigned to the (categorical) education attribute values, used to compute the difference in education level shown in Figure 4(b).
@@
Table 1: Short and full description of all races
@@
Education Numerical Value
High School Diploma 1
High school diploma 1
Associate’s degree 2
Some college 2
Bachelor’s degree 3
Master’s degree 4
Ph.D. 5
Law Degree 5
Law degree 5
Juris Doctor 5
Medical Degree 5
Medical degree 5
Dental degree 5
Dentistry degree 5

### v3

@@ 4.2 How similar is counterfactually generated text to the factual one?
Experimental setup. We first use the implementation of our model on Llama 3 8B-Instruct and Ministral-8B-Instruct to generate (factual) outputs to $2{,}000$ question prompts sourced from the LMSYS Chat 1M dataset [55]. As a system prompt we use “Keep your replies short and to the point.”. Further, for each factual output, we perform two interventions where we replace a randomly selected token $t_{i}$ with a token $t^{\prime}\neq t_{i}$. [^9] To select $t^{\prime}$, we set the probability of $t_{i}$ in $d_{i}$ to 00, re-scale the values of $d_{i}$ and use top-$p$ sampling with $p=0.9$. One of the two interventions restricts the choice of $t_{i}$ to the first half of the output sequence and the other restricts it to the second half. Then, for each intervened factual output, we feed the concatenation of the question prompt and the first part of the intervened factual output up to and including token $t^{\prime}$ as input to our model. We regenerate the second part of the output after token $t^{\prime}$ using (i) interventional token generation and (ii) counterfactual token generation, as described in Section 4.1. Finally, we measure the lexicographic similarity between the regenerated second part of the output and its factual counterpart using their (normalized) Levenshtein edit distance [56]. In our experiments, we implement our model using the Gumbel-Max SCM defined in Equation 3 as well as the top-$p$ Gumbel-Max SCM and top-$k$ Gumbel-Max SCM discussed at the end of Section 3.
Results. Figure 4 summarizes the results, which show that, across different SCMs and LLMs, the output sequences generated using counterfactual token generation are more similar to the factual sequences (i.e., the edit distance is lower) than the output sequences generated using interventional token generation. This suggests that, even though the top-$p$ and top-$k$ Gumbel-Max SCMs are not guaranteed to satisfy counterfactual stability, in practice, counterfactual token generation under both SCMs does “prioritize” selecting the same tokens $T_{i}$ that were selected during the factual generation. Although this pattern is consistent across the two LLMs, we observe that output sequences generated by both interventional and counterfactual token generation are more similar to their respective factual sequences for Llama 3 8B-Instruct than Ministral-8B-Instruct.
(a) Change in income upon intervention on sex (direct effect)
(b) Change in income upon intervention on sex (total effect)
(c) Distribution of factual
and counterfactual income
Figure 5: Comparison between factual and counterfactual income. Panel (a) shows the change in income of male (female) individuals had they been female (male), while keeping fixed the rest of their attributes preceding income in the output sequence. Panel (b) shows the change in income of male (female) individuals had they been female (male), while keeping fixed the attributes preceding sex but allowing the attributes between sex and income to change in the output sequence. Panel (c) shows the factual distributions of income of female and male individuals and the counterfactual distribution of income of female individuals under the same intervention as in panel (b). Enlarged points in panels (a, b) and dashed lines in panel (c) correspond to the median income. In all panels, we use Llama 3 8B-Instruct and set the temperature parameter to $\tau=0.8$.
@@ 4.3 Does counterfactual token generation reveal model biases?
Common approaches to addressing questions of bias and fairness rely on making counterfactual comparisons based on sensitive attributes [57]. For example, would a person’s income have been the same if their race or sex were different? In this section, we focus on a census data generation task, and demonstrate the use of counterfactual token generation to investigate potential biases of an LLM towards demographic groups.
Experimental setup. We first use the implementation of our model on Llama 3 8B-Instruct and Ministral-8B-Instruct to generate (factual) census data. For each model, we use the same input prompt three times with different seeds (see Appendix B for details), requesting $50$ individuals each time. The attributes of generated individuals include their age, sex, citizenship, race, ethnicity, marital status, number of children, occupation, income, and education, in this given order, and we discard individuals for whom the generated income is exactly zero. The factual data generated by the models result in a corpus of $114$ and $158$ fictional individuals for Llama 3 8B-Instruct and Ministral-8B-Instruct, respectively.
For each fictional person, we consider all possible interventions on each of the sensitive attributes of sex and race. For each intervention on sex and race, we concatenate the input prompt with the initial part of the output that includes the fictional person’s description up to and including the intervened sensitive attribute. This concatenated input is then used by our model to regenerate the latter part of the output, following the intervention, using counterfactual token generation (i.e., Algorithm 1). Then, we compare the factual and counterfactual values of income, education and occupation to measure the total effect of sex and race on those attributes. In addition, for each intervention on sex, we also measure the direct effect of sex on income. To this end, we concatenate the input prompt with the initial part of the output that includes the intervened value of sex and the factual values of all other attributes preceding income. This concatenated input is then used by our model to regenerate the latter part of the output starting from the income attribute. [^10] For further details regarding the difference between total and direct effects, refer to Pearl [15], Chapter 4.
(a) Change in education level
@@
Figure 6: Comparison between factual and counterfactual education and occupation. Panel (a) shows the average difference in the education level of individuals of each race had their race been different. Here, positive values indicate an improvement in education and negative values indicate a decline. Panel (b) shows the distribution shift of occupations among Asian American individuals had they been Black or African American. Green (red) sections indicate the counterfactual increase (decrease) in the number of individuals that practice each occupation. In both panels, we use Ministral-8B-Instruct and set the temperature parameter to $\tau=0.8$.
Results. Figure 5 summarizes the results with respect to the effect of sex on income using Llama 3 8B-Instruct. For both the direct and the total effect of sex on income, Figures 5(a) and  5(b) show that the generated income for most male (female) individuals would have remained unchanged or decreased (increased) had they been female (male), however, the total effect exhibits larger variance. This suggests that, in the LLM’s world model, both the direct and total effects of sex on income are present, but at a moderate level. Interestingly, the (potentially biased) effect of sex on income cannot be identified solely from the factual distributions of female and male income, which present exactly the same median, as shown in Figure 5(c). For additional results using Llama 3 8B-Instruct and Ministral-8B-Instruct, refer to Appendix C.
@@
Figure 6 summarizes the results with respect to the effect of race on education and occupation using Ministral-8B-Instruct. Figure 6(a) shows that, for individuals of all (generated) races, there exists at least one other race that, had they belonged to it, they would have experienced a significant increase or decrease in their education level (refer to Appendix B for the assignment of each education level to a numerical value and for the factual distribution of individuals across races). Specifically, we observe a consistent pattern in which the LLM would have increased the level of education for (i) American Indian or Alaska Native (Native) and (ii) Native Hawaiian or Other Pacific Islander (Hawaiian) individuals had they belonged to any other race. In contrast, it would have decreased the level of education for Asian American (Asian) individuals. [^11] The terms for races in parentheses and in Figure 6(a) are shortened descriptions, which we use for brevity. Refer to Appendix B for a list of all official descriptions. Figure 6(b) shows that, for Asian American individuals, their occupation would have shifted from STEM to humanities-related occupations had they been Black or African American. In Appendix C, we present qualitatively similar results using Llama 3 8B-Instruct.
5 Discussion and Limitations
In this section, we discuss several assumptions and limitations of our work, pointing out avenues for future research.
Methodology. Our causal model of the autoregressive process underpinning large language models operationalizes the counterfactual stability property using the Gumbel-Max SCM. It would be interesting to understand the sensitivity of counterfactual token generation to this specific choice of SCM and implement counterfactual token generation using alternative SCMs obeying counterfactual stability [58, 59]. In this context, it is also important to acknowledge that, while counterfactual stability is somewhat an appealing property, Haugh and Singal [59] have recently argued that its appropriateness depends on the application and should be justified by domain specific knowledge. Moreover, they have shown that there are cases where counterfactual stability may permit counterfactuals that it was designed to exclude. Motivated by this observation, in Appendix D, we include additional experiments on counterfactual token generation using a classical sampler for categorical distributions, which can be viewed as an SCM that does not satisfy counterfactual stability.
Further, there are reasons to believe that counterfactual statements generated using our model may be inconsistent with the underlying causal model of the world shared by humans [29, 30, 31, 32] and this may render them unreliable for certain uses cases. An interesting future direction would be to explore the use of our methodology in conjunction with human feedback to train (or fine-tune) LLMs that better understand causal relationships.
Applications. Counterfactual token generation may be useful for understanding the inner workings of an LLM. For example, it may be used as a tool for evaluating causal dependencies among physical or social attributes within the world model learned by an LLM, as in our bias experiments in Section 4.3, or quantifying the “importance” of different parts of a model’s output in reaching a final conclusion, similarly as in feature attribution [60, 61]. Further, counterfactual token generation may also be useful for creating novel interfaces for LLM-human collaborations. For example, it could be used in scenarios in which a user is satisfied with a portion of an LLM’s output but wishes to modify a few words while maintaining similarity between the subsequent text and the original output. Additionally, counterfactual token generation may be proven useful in (applications in) linguistics and cognitive science.
Evaluation. We have implemented our model on two LLMs, namely Llama 3 8B-Instruct and Ministral-8B-Instruct. It would be useful to implement our model on other LLMs and use counterfactual token generation to reveal similarities and differences between the underlying models of the world constructed by different LLMs. In this context, it would be insightful to see whether the sensitivity of an LLM’s counterfactual output changes as its number of parameters increases.
@@ Appendix A Additional counterfactual stories
In this section, we provide additional details about the census generation experiment discussed in Section 4.3.
Data generation. Figure 8 shows the complete system and user prompts used to generate the census data. For race and ethnicity, we instructed the two LLMs, through the system prompt, to select values among those reported in the latest (2020) US Census. We used this prompt three times with different seeds. Despite our request for $50$ individuals per generation, Llama 3 8B-Instruct only generated $45$, $34$ and $48$ individuals each time, resulting in a total of $127$ individuals, and Ministral-8B-Instruct generated $72$, $42$ and $44$ individuals each time, resulting in a total of $158$ individuals. The census data for Llama 3 8B-Instruct included individuals for whom the generated income is exactly zero, which we discarded, resulting in a total of $114$ individuals. Additionally, for some interventions, the models generated incomplete counterfactual outputs; we excluded those cases from further analysis. Table 3 and Figure 9 show the factual distributions of sex and race for the individuals generated by each model.
@@
Description of race values and education level. Table 1 contains the full description of each race attribute value, of which shortened versions were used in Figure 6(a). Table 2 lists the numerical values assigned to the (categorical) education attribute values, which were used to compute the difference in education level shown in Figure 6(a).
@@
Table 1: Short and full description of all races
Education Numerical Value
High school diploma, High school, 1
High School Diploma, High School
Associate’s degree, Associate’s Degree,
Associate degree, Associate Degree,
Associate’s, Associate, Undergraduate, 2
Some college, Some College, College,
Vocational Training
Bachelor’s degree, Bachelor’s Degree, 3
Bachelor’s, Nursing Degree
Master’s degree, Master’s Degree, 4
Master’s
Ph.D., PhD, Doctorate degree,
Doctorate Degree, Doctorate,
Doctoral degree, Doctoral Degree,
JD, Juris Doctor, Juris Doctor (JD), 5
Law degree, Law degree, PharmD,
Pharmacy Degree, Dental degree,
Dental Degree, Dentistry degree,
MD, Medical degree, Medical Degree
@@
Model Male Female
Llama 3 8B-Instruct 72 55
Ministral-8B-Instruct 79 79
Table 3: Number of male and female individuals generated by each model
Figure 9: Factual distributions of race. The panels show the empirical factual distributions of race for the individuals generated by Llama 3 8B-Instruct (left) and Ministral-8B-Instruct (right).
Appendix C Additional experimental results on bias detection
Figure 10: Comparison between factual and counterfactual income. The left (right) panel shows the factual distributions of income of female (male) individuals and their counterfactual distribution of income under two interventions: (i) had they been male (female), while keeping fixed the rest of their attributes preceding income in the output sequence, and (ii) had they been male (female), while keeping fixed the attributes preceding sex but allowing the attributes between sex and income to change in the output sequence. In both panels, dashed lines correspond to the median income. Here, we use Llama 3 8B-Instruct and set the temperature parameter to $\tau=0.8$.
(a) Change in income upon intervention on sex (direct effect)
(b) Change in income upon intervention on sex (total effect)
(c) Distributions of factual and counterfactual income
Figure 11: Comparison between factual and counterfactual income. Panel (a) shows the change in income of male (female) individuals had they been female (male), while keeping fixed the rest of their attributes preceding income in the output sequence. Panel (b) shows the change in income of male (female) individuals had they been female (male), while keeping fixed the attributes preceding sex but allowing the attributes between sex and income to change in the output sequence. Panel (c) shows the factual distributions of income of female and male individuals and their counterfactual distributions of income under the same interventions as in panels (a) and (b). Enlarged points in panels (a, b) and dashed lines in panel (c) correspond to the median income. In all panels, we use Ministral-8B-Instruct and set the temperature parameter to $\tau=0.8$.
Here, we present experimental results complementing those in Section 4.3, regarding the direct and total effects of sex on income and the total effects of race on education and occupation. Figure 10 complements Figure 5(c); it shows the distributions of the factual and counterfactual incomes of the individuals generated by Llama 3 8B-Instruct, under all possible interventions on their sex. Figure 11 complements Figure 5; it summarizes the results regarding direct and total effects of sex on income for the individuals generated by Ministral-8B-Instruct. Figure 12 complements Figure 6; it summarizes the results regarding total effects of race on education and occupation for the individuals generated by Llama 3 8B-Instruct.
(a) Change in education level upon intervention on race
(b) Change in occupation upon intervention on race
Figure 12: Comparison between factual and counterfactual education and occupation. Panel (a) shows the average difference in the education level of individuals of each race had their race been different. Here, positive values indicate an improvement in education and negative values indicate a decline. Panel (b) shows the distribution shift of occupations among Asian American individuals had they been Black or African American. Green (red) sections indicate the counterfactual increase (decrease) in the number of individuals that practice each occupation. In both panels, we use Llama 3 8B-Instruct and set the temperature parameter to $\tau=0.8$.
Appendix D Additional experiments using a sampler that does not satisfy counterfactual stability
In this section, we conduct counterfactual token generation using a classical sampler for categorical distributions, which can be viewed as an SCM that is not guaranteed to satisfy counterfactual stability.
Experimental setup. We experiment with two different SCMs to implement the function $f_{T}$ responsible for sampling the next token on Llama 3 8B-Instruct: (i) the Gumbel-Max SCM defined in Eq. 3 and (ii) an SCM based on inverse transform sampling defined as follows:
$f_{T}(D_{i},U_{i})=\mathop{\mathrm{argmin}}_{j\in\{1,2,\ldots,|V|\}}\left\{j\cdot\mathds{1}\left[\sum_{k=1}^{j}D_{i,k}\geq U_{i}\right]\right\}.$ (4)
Under the SCM based on inverse transform sampling, the next token is sampled as follows. Let the tokens $t\in V$ follow a fixed order across time steps $i$. To generate a token $T_{i}$, for each position $j\in\{1,2,\ldots,|V|\}$, the SCM based on inverse transform sampling computes the cumulative sum of probabilities in the distribution $D_{i}$ corresponding to the tokens in positions $k\leq j$. Then, it samples a unidimensional noise variable $U_{i}\sim\text{Uniform}(0,1)$ and selects the first token in the vocabulary $V$ whose corresponding cumulative sum is greater than or equal to $U_{i}$.
Similarly as in Section 4.2, we compare the edit distance between factual output sequences and sequences generated through interventional and counterfactual token generation using the Gumbel-Max SCM and the SCM given by Eq. 4.
Results. Figure 13 summarizes the results, which show that, under both SCMs, the output sequences generated using counterfactual token generation remain more similar to their respective factual sequence (i.e., exhibit lower edit distance) compared to the sequences generated using interventional token generation. Moreover, as expected, the counterfactual output sequences generated using the SCM defined by Eq. 4, which does not satisfy the property of counterfactual stability, exhibit an edit distance higher than those generated using the Gumbel-Max SCM.
Figure 13: Comparison between interventional and counterfactual token generation under two different SCMs. The plot shows the edit distance between the factual sequence of tokens and the sequence generated by interventional and counterfactual token generation using the Gumbel-Max SCM (GM) defined by Eq. 3 and the SCM based on inverse transform sampling (ITS) defined by Eq. 4, against various values of the temperature parameter $\tau$. The edit distance is averaged over $4{,}000$ output sequences, resulting from two independent interventions per factual sequence, and error bars represent $95\%$ confidence intervals. In this panel, we use Llama 3 8B-Instruct.